OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#9398 — Lille optical loop
Scheduled Maintenance Report for Web Cloud
Completed
Thereafter an optical cutoff on the 23th September. An optical maintenance is expected on the cable to Lille (see FS#5477).

From 10:00pm a team will intervene in order to repair the cable completely, the fiber will be cutoff during the reparation. The security in Lille will be impacted and the NRA in Canteleu CAN59 and Haubourdin HAU59 will remain unavailable.



Update(s):

Date: 2013-09-24 22:49:58 UTC
The security is back to normal. The NRA in Canteleu CAN59 and Haubourdin HAU59 are no longer isolated.
Posted Sep 24, 2013 - 22:48 UTC